# Communication-Efficient Module-Wise Federated Learning for Grasp Pose Detection in Cluttered Environments

### v1

@@ Abstract
Grasp pose detection (GPD) is a fundamental capability for robotic autonomy, but its reliance on large, diverse datasets creates significant data privacy and centralization challenges. Federated Learning (FL) offers a privacy-preserving solution, but its application to GPD is hindered by the substantial communication overhead of large models, a key issue for resource-constrained robots. To address this, we propose a novel module-wise FL framework that begins by analyzing the learning dynamics of the GPD model’s functional components. This analysis identifies slower-converging modules, to which our framework then allocates additional communication effort. This is realized through a two-phase process: a standard full-model training phase is followed by a communication-efficient phase where only the identified subset of slower-converging modules is trained and their partial updates are aggregated. Extensive experiments on the GraspNet-1B dataset demonstrate that our method outperforms standard FedAvg and other baselines, achieving higher accuracy for a given communication budget. Furthermore, real-world experiments on a physical robot validate our approach, showing a superior grasp success rate compared to baseline methods in cluttered scenes. Our work presents a communication-efficient framework for training robust, generalized GPD models in a decentralized manner, effectively improving the trade-off between communication cost and model performance.
@@ I Introduction
Grasp pose detection (GPD) is a fundamental capability for autonomous robotic systems, with applications ranging from industrial automation to assistive robotics. In recent years, learning-based approaches have become the dominant paradigm for this task [1, 2]. This success, however, is critically dependent on large-scale, diverse datasets capable of fostering generalization across various objects, scenes, and environmental conditions [3, 4, 2, 5].
However, acquiring such data presents substantial challenges. The collection process is not only laborious but, more critically, raises significant privacy concerns when data originates from private settings like homes or proprietary industrial environments [6, 2, 7]. The traditional requirement of centralizing sensitive sensor information thus becomes a major obstacle, restricting the ability to pool datasets from multiple sources. This limitation ultimately hinders the training of more robust, generalized models and creates a data island problem that impedes the development of universally applicable GPD.
@@
Federated Learning (FL) has emerged as a compelling paradigm for addressing privacy and data governance challenges, enabling multiple clients to collaboratively train a model without exchanging their raw data [7]. This decentralized approach is highly pertinent to robotics, allowing for the use of diverse, distributed datasets for applications like autonomous navigation [8] and manipulator control [9] while preserving data privacy. However, a significant bottleneck in FL is the communication overhead, particularly for the large-parameter models used in GPD [7]. The frequent transmission of model updates consumes substantial bandwidth and energy, a critical issue for resource-constrained robots. Although prior works like [10] have aimed to address data heterogeneity by clustering robots (or clients), their approach can intensify this resource burden.
To address the resource burden of FL for GPD, we observe that contemporary models in this domain feature multi-modular architectures [2, 11, 12, 13]. To analyze these architectures in an FL setting, we measured the cosine similarity of model updates on a module-by-module basis [14, 15]. Our analysis reveals that these modules exhibit heterogeneous learning dynamics, allowing us to distinguish between faster-converging modules and the slower-converging ones that act as training bottlenecks. This finding motivates our proposal of a novel, two-phase module-wise FL algorithm, as illustrated in Figure 1. The first phase consists of standard full-model training and full aggregation. The second phase, in contrast, employs a parameter-efficient approach by strategically concentrating resources: only the subset of modules identified as slower-converging undergoes partial-model training, and subsequently, only their updated parameters are aggregated. This approach maximizes the performance return on a given communication budget, enhancing the feasibility of FL for complex robotics tasks like GPD. Our primary contributions are as follows:
@@
  • Motivated by our analysis, we propose a novel module-wise FL algorithm featuring a two-phase process. This algorithm strategically allocates additional training and communication resources exclusively to the identified slower-converging modules, enhancing communication efficiency and maximizing performance for a given budget.
  • Through extensive experiments on the GraspNet-1B dataset [2] and in the real world, we demonstrate the superiority of our approach over existing baselines and validate its practical effectiveness.
@@ II-A Data-Driven Grasp Pose Detection
GPD aims to identify a stable 6-DoF (Degree of freedom) gripper pose from sensory inputs like RGB-D images or 3D point clouds. Historically, this problem was often addressed through model-based approaches [1], which relied on having precise 3D models of the objects. However, these methods faced significant limitations, such as difficulty handling novel objects and sensitivity to sensor noise. The advent of deep learning has since spurred a paradigm shift towards data-driven grasping methodologies that learn directly from data. The field’s progress is driven by advanced datasets [3, 2, 5] and novel model architectures [16, 17, 12]. Early data-driven works focused on representing grasps as 2D rectangles in the image plane [18, 16, 4]. However, these 2D approaches could not represent the full spatial orientation required for manipulation in cluttered scenes. This limitation spurred a paradigm shift towards methods that detect full 6-DoF poses directly from 3D point clouds. These modern approaches have progressed from sampling-based techniques that evaluate numerous candidates [2] to more efficient, single-stage frameworks that directly predict grasp poses [17, 19, 20, 12]. This advancement is strongly supported by the development of large-scale datasets, which are crucial for training and evaluation. These range from early synthetic datasets with millions of grasp examples [3] to recent real-world benchmarks focusing on cluttered environments and object diversity [2, 5]. However, the very success built upon these diverse, real-world datasets underscores a critical challenge: the logistical and privacy-related difficulties in centralizing such large-scale, sensitive data.
@@ II-B Federated Learning for Robotic Grasping
FL has emerged as a machine learning paradigm that enables multiple decentralized clients to collaboratively train a model without sharing their raw local datasets [7]. In this framework, exemplified by the canonical FedAvg algorithm, clients locally train a shared model, and a central server aggregates only their resulting updates to produce an improved global model. The inherent privacy preservation of this framework makes it highly suitable for robotics, where applications often involve sensitive data from private residences or proprietary industrial processes. Reflecting its growing importance, FL is now being applied to a range of robotic tasks, including multi-robot trajectory prediction [21], visual robotic navigation [8], and robotic manipulation [9].
However, applying FL to GPD is still in its early stages. While a prior work by [10] did explore this direction by applying a clustering-based FL approach to 2D grasping, it suffers from two critical limitations. First, from a task perspective, it is confined to simplified 2D rectangular grasps, which are insufficient for dexterous manipulation in cluttered 3D scenes. Second, from a methodological standpoint, its clustering process not only incurs significant communication overhead but also risks harming generalization by permanently excluding clients with unique data distributions. Therefore, applying FL to the more practical and challenging GPD problem, while simultaneously addressing the communication-efficiency of the FL algorithm itself, remains a significant open challenge. Our work aims to fill this critical gap by proposing a novel, communication-efficient FL strategy specifically designed for multi-modular GPD models.
@@ II-C Communication-Efficient Federated Learning
A primary bottleneck in FL is the substantial communication overhead required for the iterative transmission of large-scale models between clients and a central server. This challenge is exacerbated as model complexity and the number of participating clients increase, creating a critical trade-off between model performance and communication cost. Consequently, a significant body of research has focused on developing communication-efficient FL strategies. Early approaches centered on reducing communication frequency by allowing for more local computation on client devices before aggregation [7]. Another major line of work involves compressing the weight updates themselves through techniques like quantization [22, 23], which reduces the precision of weight updates, and sparsification [24], which transmits only a subset of the most significant updates. More recently, sophisticated adaptive methods have been introduced. An example is FedLAMA [25], which employs a layer-wise adaptive aggregation strategy. Motivated by the observation that different layers converge at different rates, this approach adjusts the aggregation frequency of each layer based on its contribution to model discrepancy. However, these methods fundamentally assume a monolithic model architecture, and consequently, their approach to efficiency is restricted to the layer level. In contrast, our work is motivated by the functionally distinct, modular structures inherent in advanced robotics models. Instead of viewing the model as a simple stack of layers, we leverage the heterogeneous learning dynamics of its functional components. This allows us to propose a novel communication-efficient strategy tailored to multi-component systems, enabling a more granular and effective allocation of communication resources.
Figure 2: Examples of multi-module architectures for grasp pose detection, from [2] (left) and [13] (right). These diagrams are simplified to show only learnable neural network modules, omitting other operations. The black arrows represent the main forward pass, though additional residual connections may exist in the actual models.
@@ III Preliminaries and Framework Overview
The weight $k_{c}=|D_{c}|/|D|$ (where $|D|=\sum_{i=1}^{n}|D_{i}|$) typically represents the relative size of each client’s dataset. The canonical algorithm to solve this objective is FedAvg [7]. The training process operates in communication rounds (indexed by $r$), where the server first distributes the current global model $w^{r}$ to all clients. Each client then performs local training to compute an updated model $w_{c}^{r+1}$ and, to preserve privacy, transmits only the model update $\Delta w_{c}^{r+1}=w_{c}^{r+1}-w^{r}$ back to the server. The server aggregates these updates to form the next global model $w^{r+1}$:
@@
While more advanced optimizers exist [26, 27], we use FedAvg [7] as it provides a clear baseline for validating our module-wise strategy, which is orthogonal to the choice of optimizer.
Grasp Pose Representation. Directly learning the 6-DoF grasp pose matrix is challenging for neural networks due to the complex constraints of rotation matrices. Consequently, many state-of-the-art approaches [2, 11, 13] adopt a decoupled representation. Following the methodology in [13], we define a 6-DoF grasp $\mathbf{G}$ as a tuple $[\mathbf{c},v,a,d,w,s]$. In this formulation, $\mathbf{c}\in\mathbb{R}^{3}$ is the grasp center point. The integers $v$, $a$, and $d$ represent the discretized approach direction, in-plane rotation, and grasp depth, respectively. Finally, $w\in\mathbb{R}$ is the grasp width and $s\in\mathbb{R}$ is the predicted grasp quality score.
@@ IV-A Analysis of Module-wise Learning Dynamics in FL
Modern deep learning models for GPD often feature multi-modular architectures [2, 11, 12, 13], as illustrated in Figure 2. For instance, the model presented in [2] consists of PointNet++ [28], ApproachNet, OperationNet, and ToleranceNet. Similarly, the architecture in [13] is composed of 3D-UNet, GraspableNet, ViewNet, CGGI-Net, and GHLI-Net. These modules are tasked with specific functions, such as extracting 3D features, determining approaching vectors, and estimating grasp quality. Motivated by this observation, we investigate whether these functionally distinct modules exhibit heterogeneous learning dynamics during FL training.
To quantify these dynamics, we analyze the alignment between local client updates and the aggregated global update for each module using cosine similarity. This metric has been employed in clustered FL research [14, 15] to assess biases in client data. For each module $m\in M$, where $M$ is the set of all modules in the model, in communication round $r$, we compute the average module-wise similarity $S_{m}^{r}$ as:
@@
Our analysis, conducted by applying FedAvg to a GPD model with a multi-modular architecture analogous to that in [13], reveals heterogeneous learning dynamics at the module level. As illustrated in Figure 3 (see Section V for experimental details), these modules exhibit distinct similarity trajectories. For example, the module analogous to GraspableNet consistently shows high cosine similarity that diminishes gradually. Conversely, the module analogous to 3D-UNet shows a lower initial similarity and a steady decline. These similarity values offer insight into the module’s convergence state, showing the difference between learning general and personalized features, a key idea in [29, 30]. Specifically, if a module is far from its stationary point, it is primarily learning general features, leading to high similarity between local and global updates. As a module approaches convergence, however, it begins to learn local-data-specific features, causing the local updates to diverge and thus lowering their similarity to the global update. A low similarity suggests that the module is close to a stationary point, and thus can be considered fast-converging. In contrast, a module that maintains a high similarity value is interpreted as being relatively far from its stationary point, signifying a slow-converging module. This leads to our core design principle: to maximize the performance return on a given communication budget, additional training and aggregation efforts should be directed towards the slower-converging modules, as they are the primary bottlenecks to overall model performance.
Input: Initial global model parameters $w^{0}$; Total communication rounds $R$; Number of clients $n$; Client weights $\{k_{c}\}_{c=1}^{n}$; Set of all modules $M$ and a subset $\tilde{M}\subset M$; Local epochs $\tau_{1},\tau_{2}$
@@ 6      $\Delta w_{g}\leftarrow\sum_{c=1}^{n}k_{c}\Delta w_{c}$;
8       Server distributes $w^{r+0.5}$ to all clients;

### v2

@@ Abstract
Grasp pose detection (GPD) is a fundamental capability for robotic autonomy, but its reliance on large, diverse datasets creates significant data privacy and centralization challenges. Federated Learning (FL) offers a privacy-preserving solution, but its application to GPD is hindered by the substantial communication overhead of large models, a key issue for resource-constrained robots. To address this, we propose a novel module-wise FL framework that begins by analyzing the learning dynamics of the GPD model’s functional components. This analysis identifies slower-converging modules, to which our framework then allocates additional communication effort. This is realized through a two-phase process: a standard full-model training phase is followed by a communication-efficient phase where only an adaptively identified subset of slower-converging modules is trained and their partial updates are aggregated. Extensive experiments on the GraspNet-1B dataset demonstrate that our method outperforms standard FedAvg and other baselines, achieving higher accuracy for a given communication budget. Furthermore, real-world experiments on a physical robot validate our approach, showing a superior grasp success rate compared to baseline methods in cluttered scenes. Our work presents a communication-efficient framework for training robust, generalized GPD models in a decentralized manner, effectively improving the trade-off between communication cost and model performance.
@@ I Introduction
Grasp pose detection (GPD) is a fundamental capability for autonomous robotic systems, with applications ranging from industrial automation to assistive robotics. In recent years, learning-based approaches have become the dominant paradigm for this task [bohg2013data, fang2020graspnet]. This success, however, is critically dependent on large-scale, diverse datasets capable of fostering generalization across various objects, scenes, and environmental conditions [mahler2016dex, levine2018learning, fang2020graspnet, vuong2024grasp].
However, acquiring such data presents substantial challenges. The collection process is not only laborious but, more critically, raises significant privacy concerns when data originates from private settings like homes or proprietary industrial environments [pinto2016supersizing, fang2020graspnet, mcmahan2017communication]. The traditional requirement of centralizing sensitive sensor information thus becomes a major obstacle, restricting the ability to pool datasets from multiple sources. This limitation ultimately hinders the training of more robust, generalized models and creates a data island problem that impedes the development of universally applicable GPD.
@@
Federated Learning (FL) has emerged as a compelling paradigm for addressing privacy and data governance challenges, enabling multiple clients to collaboratively train a model without exchanging their raw data [mcmahan2017communication]. This decentralized approach is highly pertinent to robotics, allowing for the use of diverse, distributed datasets for applications like autonomous navigation [gummadi2024fed] and manipulator control [wang2024grasp] while preserving data privacy. However, a significant bottleneck in FL is the communication overhead, particularly for the large-parameter models used in GPD [mcmahan2017communication]. The frequent transmission of model updates consumes substantial bandwidth and energy, a critical issue for resource-constrained robots. Although prior works like [kang2023fogl] have aimed to address data heterogeneity by clustering robots (or clients), their approach can intensify this resource burden.
To address the resource burden of FL for GPD, we observe that contemporary models in this domain feature multi-modular architectures [fang2020graspnet, wang2021graspness, wang2023granet, wu2024economic]. To analyze these architectures in an FL setting, we measured the cosine similarity of model updates on a module-by-module basis [sattler2020clustered, ma2022convergence]. Our analysis reveals that these modules exhibit heterogeneous learning dynamics, allowing us to distinguish between faster-converging modules and the slower-converging ones that act as training bottlenecks. This finding motivates our proposal of a novel, two-phase module-wise FL algorithm, as illustrated in Figure 1. The first phase consists of standard full-model training and full aggregation. The second phase, in contrast, employs a parameter-efficient approach by strategically concentrating resources: only the subset of modules identified as slower-converging undergoes partial-model training, and subsequently, only their updated parameters are aggregated. This approach maximizes the performance return on a given communication budget, enhancing the feasibility of FL for complex robotics tasks like GPD. Our primary contributions are as follows:
@@
  • Motivated by our analysis, we propose a novel module-wise FL algorithm featuring a two-phase process. This algorithm allocates additional training and communication resources exclusively to the slower-converging modules adaptively, enhancing communication efficiency and maximizing performance for a given budget.
  • Through extensive experiments on the GraspNet-1B dataset [fang2020graspnet] and in the real world, we demonstrate the superiority of our approach over existing baselines and validate its practical effectiveness.
@@ II-A Data-Driven Grasp Pose Detection
GPD aims to identify a stable 6-DoF (Degree of freedom) gripper pose from sensory inputs like RGB-D images or 3D point clouds. Historically, this problem was often addressed through model-based approaches [bohg2013data], which relied on having precise 3D models of the objects. However, these methods faced significant limitations, such as difficulty handling novel objects and sensitivity to sensor noise. The advent of deep learning has since spurred a paradigm shift towards data-driven grasping methodologies that learn directly from data. The field’s progress is driven by advanced datasets [mahler2016dex, fang2020graspnet, vuong2024grasp] and novel model architectures [zhang2019roi, wang2023granet]. Early data-driven works focused on representing grasps as 2D rectangles in the image plane [lenz2015deep, levine2018learning]. However, these 2D approaches could not represent the full spatial orientation required for manipulation in cluttered scenes. This limitation spurred a paradigm shift towards methods that detect full 6-DoF poses directly from 3D point clouds. These modern approaches have progressed from sampling-based techniques that evaluate numerous candidates [fang2020graspnet] to more efficient, single-stage frameworks that directly predict grasp poses [zhang2019roi, liu2022transgrasp, ma2023towards, wang2023granet]. This advancement is strongly supported by the development of large-scale datasets, which are crucial for training and evaluation. These range from early synthetic datasets with millions of grasp examples [mahler2016dex] to recent real-world benchmarks focusing on cluttered environments and object diversity [fang2020graspnet, vuong2024grasp]. However, the very success built upon these diverse, real-world datasets underscores a critical challenge: the logistical and privacy-related difficulties in centralizing such large-scale, sensitive data.
@@ II-B Federated Learning for Robotic Grasping
FL has emerged as a machine learning paradigm that enables multiple decentralized clients to collaboratively train a model without sharing their raw local datasets [mcmahan2017communication]. In this framework, exemplified by the canonical FedAvg algorithm, clients locally train a shared model, and a central server aggregates only their resulting updates to produce an improved global model. The inherent privacy preservation of this framework makes it highly suitable for robotics, where applications often involve sensitive data from private residences or proprietary industrial processes. Reflecting its growing importance, FL is now being applied to a range of robotic tasks to scaling robot learning. It includes multi-robot trajectory prediction [majcherczyk2021flow], visual robotic navigation [gummadi2024fed], and robotic manipulation [wang2024grasp], [miao2025fedvla].
However, applying FL to GPD is still in its early stages. While a prior work by [kang2023fogl] did explore this direction by applying a clustering-based FL approach to 2D grasping, it suffers from two critical limitations. First, from a task perspective, it is confined to simplified 2D rectangular grasps, which are insufficient for dexterous manipulation in cluttered 3D scenes. Second, from a methodological standpoint, its clustering process not only incurs significant communication overhead but also risks harming generalization by permanently excluding clients with unique data distributions. Therefore, applying FL to the more practical and challenging GPD problem, while simultaneously addressing the communication-efficiency of the FL algorithm itself, remains a significant open challenge. Our work aims to fill this critical gap by proposing a novel, communication-efficient FL strategy specifically designed for multi-modular GPD models.
@@ II-C Communication-Efficient Federated Learning
A primary bottleneck in FL is the substantial communication overhead required for the iterative transmission of large-scale models between clients and a central server. This challenge is exacerbated as model complexity and the number of participating clients increase, creating a critical trade-off between model performance and communication cost. Consequently, a significant body of research has focused on developing communication-efficient FL strategies. Early approaches centered on reducing communication frequency by allowing for more local computation on client devices before aggregation [mcmahan2017communication]. Another major line of work involves compressing the weight updates themselves through techniques like quantization [konevcny2016federated, alistarh2017qsgd], which reduces the precision of weight updates, and sparsification [sattler2019robust], which transmits only a subset of the most significant updates. More recently, sophisticated adaptive methods have been introduced. An example is FedLAMA [lee2023layer], which employs a layer-wise adaptive aggregation strategy. Motivated by the observation that different layers converge at different rates, this approach adjusts the aggregation frequency of each layer based on its contribution to model discrepancy. However, these methods fundamentally assume a monolithic model architecture, and consequently, their approach to efficiency is restricted to the layer level. In contrast, our work is motivated by the functionally distinct, modular structures inherent in advanced robotics models. Instead of viewing the model as a simple stack of layers, we leverage the heterogeneous learning dynamics of its functional components. This allows us to propose a novel communication-efficient strategy tailored to multi-component systems, enabling a more granular and effective allocation of communication resources.
Figure 2: Examples of multi-module architectures for GPD, from [fang2020graspnet] (left) and [wu2024economic] (right). These diagrams are simplified to show only learnable neural network modules, omitting other operations. The black arrows represent the main forward pass, though additional residual connections may exist in the actual models.
@@ III Preliminaries and Framework Overview
The weight $k_{c}=|D_{c}|/|D|$ (where $|D|=\sum_{i=1}^{n}|D_{i}|$) typically represents the relative size of each client’s dataset. The canonical algorithm to solve this objective is FedAvg [mcmahan2017communication]. The training process operates in communication rounds (indexed by $r$), where the server first distributes the current global model $w^{r}$ to all clients. Each client then performs local training to compute an updated model $w_{c}^{r+1}$ and, to preserve privacy, transmits only the model update $\Delta w_{c}^{r+1}=w_{c}^{r+1}-w^{r}$ back to the server. The server aggregates these updates to form the next global model $w^{r+1}$:
@@
While more advanced optimizers exist [li2020federated, karimireddy2020scaffold], we use FedAvg [mcmahan2017communication] as it provides a clear baseline for validating our module-wise strategy, which is orthogonal to the choice of optimizer.
Grasp Pose Representation. Directly learning the 6-DoF grasp pose matrix is challenging for neural networks due to the complex constraints of rotation matrices. Consequently, many state-of-the-art approaches [fang2020graspnet, wang2021graspness, wu2024economic] adopt a decoupled representation. Following the methodology in [wu2024economic], we define a 6-DoF grasp $\mathbf{G}$ as a tuple $\mathbf{c},v,a,d,w,s]$. In this formulation, $\mathbf{c}\in\mathbb{R}^{3}$ is the grasp center point. The integers $v$, $a$, and $d$ represent the discretized approach direction, in-plane rotation, and grasp depth, respectively. Finally, $w\in\mathbb{R}$ is the grasp width and $s\in\mathbb{R}$ is the predicted grasp quality score.
@@ IV-A Analysis of Module-wise Learning Dynamics in FL
Modern deep learning models for GPD often feature multi-modular architectures [fang2020graspnet, wang2021graspness, wang2023granet, wu2024economic], as illustrated in Figure 2. This modularity is a notable characteristic of robotics models, stemming from the inherent difficulty of training a neural network to interpret complex 3D spatial information [fang2020graspnet]. Consequently, a common architectural approach is to employ a combination of multiple specialized modules. For instance, the model presented in [fang2020graspnet] consists of PointNet++ [qi2017pointnet++], ApproachNet, OperationNet, and ToleranceNet. Similarly, the architecture in [wu2024economic] is composed of 3D-UNet, GraspableNet, ViewNet, CGGI-Net, and GHLI-Net. These modules are tasked with specific functions, such as extracting 3D features, determining approaching vectors, and estimating grasp quality. Motivated by this observation, we investigate whether these functionally distinct modules exhibit heterogeneous learning dynamics during FL training.
To quantify these dynamics, we analyze the alignment between local client updates and the aggregated global update for each module using cosine similarity. This metric has been employed in clustered FL research [sattler2020clustered, ma2022convergence] to assess biases in client data. For each module $m\in M$, where $M$ is the set of all modules in the model, in communication round $r$, we compute the average module-wise similarity $S_{m}^{r}$ as:
@@
Our analysis, conducted by applying FedAvg to a GPD model with a multi-modular architecture analogous to that in [wu2024economic], reveals heterogeneous learning dynamics at the module level. As illustrated in Figure 3 (see Section V for experimental details), these modules exhibit distinct similarity trajectories. For example, the module analogous to GraspableNet consistently shows high cosine similarity that diminishes gradually. Conversely, the module analogous to 3D-UNet shows a lower initial similarity and a steady decline. These similarity values offer insight into the module’s convergence state, showing the difference between learning general and personalized features, a key idea in [wang2022personalized, yi2024pfedmoe]. Specifically, if a module is far from its stationary point, it is primarily learning general features, leading to high similarity between local and global updates. As a module approaches convergence, however, it begins to learn local-data-specific features, causing the local updates to diverge and thus lowering their similarity to the global update. A low similarity suggests that the module is close to a stationary point, and thus can be considered fast-converging. In contrast, a module that maintains a high similarity value is interpreted as being relatively far from its stationary point, signifying a slow-converging module. This leads to our core design principle: to maximize the performance return on a given communication budget, additional training and aggregation efforts should be directed towards the slower-converging modules, as they are the primary bottlenecks to overall model performance.
Input: Initial global model parameters $w^{0}$; Total communication rounds $R$; Number of clients $n$; Client weights $\{k_{c}\}_{c=1}^{n}$; Local epochs $\tau_{1},\tau_{2}$; Number of slow-converging modules to select $\zeta$
@@ 6 $\Delta w_{g}\leftarrow\sum_{c=1}^{n}k_{c}\Delta w_{c}$;
8  Calculate $S_{m}^{r}$ using updates from Phase 1;
9  $\tilde{M}^{r}\leftarrow$ Select top-$\zeta$ modules with highest $S_{m}^{r}$;
10   Server distributes $w^{r+0.5}$ and $\tilde{M}^{r}$ to all clients;